# Stability and Decay for the 2D Anisotropic Navier-Stokes Equations with Fractional Horizontal Dissipation on $\mathbb{R}^{2}$

### v1

@@ 2 Anisotropic Estimates and Tools
If $\|\Lambda_{1}^{-\sigma}u_{0}\|_{H^{k}}<\varepsilon$ with $\varepsilon<\nu/C$, then a standard bootstrap argument applied to (4.18) yields
$\|\Lambda_{1}^{-\sigma}u(t)\|_{H^{k}}^{2}+\nu\int_{0}^{t}\|\Lambda_{1}^{s}\Lambda_{1}^{-\sigma}u(\tau)\|_{H^{k}}^{2}\,d\tau\leq C\|\Lambda_{1}^{-\sigma}u_{0}\|_{H^{k}}^{2}\leq C\varepsilon^{2}.$
We now turn to the proof of Lemma 4.1.
Proof of Lemma 4.1. We decompose $J$ into two parts:
$J$ $=\sum_{\beta=1}^{k}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{\ell}^{\beta}u\,\partial_{1}\partial_{\ell}^{k-\beta}u\big)\,\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\,dx$
$=\sum_{\beta=1}^{k-1}\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{\ell}^{\beta}u\,\partial_{1}\partial_{\ell}^{k-\beta}u\big)\,\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\,dx+\int_{\mathbb{R}^{2}}\Lambda_{1}^{-\sigma}\big(\partial_{\ell}^{k}u\,\partial_{1}u\big)\,\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\,dx$
$=:J_{1}+J_{2}.$
Since $s<\tfrac{1}{2}+\sigma$, we may employ Cauchy-Schwarz inequality and then apply Lemma 2.3 to estimate $J_{1}$:
$J_{1}$ $\leq\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{-\sigma}(\partial_{\ell}^{\beta}u\,\partial_{1}\partial_{\ell}^{k-\beta}u)\|_{L^{2}}\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}$
$\leq\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{-\sigma+s}\partial_{\ell}^{\beta}u\|_{L_{x_{1}}^{2}L_{x_{2}}^{\infty}}\,\|\Lambda_{1}^{\frac{1}{2}-s}\partial_{1}\partial_{\ell}^{k-\beta}u\|_{L^{2}}\,\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}$
$\leq\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{-\sigma+s}\partial_{\ell}^{\beta}u\|_{L_{x_{1}}^{2}H_{x_{2}}^{1}}\,\|\Lambda_{1}^{-\sigma+s}\Lambda_{1}^{\frac{3}{2}-2s+\sigma}\partial_{\ell}^{k-\beta}u\|_{L^{2}}\,\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}.$
By the admissible ranges of $\beta$, $s$, and $\sigma$, we have
$\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{-\sigma+s}\partial_{\ell}^{\beta}u\|_{L_{x_{1}}^{2}H_{x_{2}}^{1}}\leq\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}},\qquad\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{-\sigma+s}\Lambda_{1}^{\frac{3}{2}-2s+\sigma}\partial_{\ell}^{k-\beta}u\|_{L^{2}}\leq\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}}.$
Consequently,
$J_{1}\leq C\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u\|_{H^{k}}.$ (4.19)
The term $J_{2}$ can be estimated in a similar manner. By Hölder’s inequality and Lemma 2.3, we obtain
$J_{2}$ $\leq\|\Lambda_{1}^{-\sigma}(\partial_{\ell}^{k}u\,\partial_{1}u)\|_{L^{2}}\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}$
$\leq\|\Lambda_{1}^{-\sigma+s}\partial_{\ell}^{k}u\|_{L^{2}}\,\|\Lambda_{1}^{\frac{1}{2}-s}\partial_{1}u\|_{L_{x_{1}}^{2}L_{x_{2}}^{\infty}}\,\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}$
$\leq\|\Lambda_{1}^{-\sigma+s}\partial_{\ell}^{k}u\|_{L^{2}}\,\|\Lambda_{1}^{-\sigma+s}\Lambda_{1}^{\frac{3}{2}-2s+\sigma}u\|_{L_{x_{1}}^{2}H_{x_{2}}^{1}}\,\|\Lambda_{1}^{-\sigma}\partial_{\ell}^{k}u\|_{L^{2}}.$ (4.20)
Combining (4.19) and (4.20), we conclude that
$J\leq C\|\Lambda_{1}^{-\sigma+s}u\|_{H^{k}}^{2}\|\Lambda_{1}^{-\sigma}u\|_{H^{k}}.$
This completes the proof of Lemma 4.1.
@@ 4.3 The decay of $u_{1}$
$\|u_{1}(t)\|_{L^{2}}$ $\leq\|e^{\nu\Lambda_{1}^{2s}t}u_{0,1}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}(u\cdot\nabla u_{1})(\tau)\|_{L^{2}}\,d\tau$
@@
$\|e^{\nu\Lambda_{1}^{2s}t}u_{0,1}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma}{2s}}\Big(\|\Lambda_{1}^{-\sigma}u_{0,1}\|_{L^{2}}+\|u_{0}\|_{L^{2}}\Big)$
@@ 4.4 The decay of $u_{2}$
$u_{2}(t)=e^{\nu\Lambda_{1}^{2s}t}u_{0,2}-\int_{0}^{t}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)d\tau.$
@@
$\|u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}u_{0,2}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}\,d\tau.$
@@
$\|e^{\nu\Lambda_{1}^{2s}t}u_{0,2}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\big(\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|u_{0,2}\|_{L^{2}}\big)$
@@
For $II_{1}$, by Lemma 2.7, we take out $\Lambda_{1}^{\sigma+\frac{2}{3}}$ and applying Lemma 2.3 and Lemma 2.4, we have
@@
$\leq$ $\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}(\Lambda_{1}^{\frac{5}{6}}u_{1})u\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}u_{1}(\Lambda_{1}^{\frac{5}{6}}u)\|_{L^{2}}\Big)d\tau.$ (4.23)
@@
$\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}(\Lambda_{1}^{\frac{5}{6}}u_{1})u\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}u_{1}(\Lambda_{1}^{\frac{5}{6}}u)\|_{L^{2}}\Big)d\tau$
@@ 4.5 The decay of $\partial_{2}u_{1}$
$\|\partial_{2}u_{1}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{2}u_{0,1}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{2}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
@@
$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{2}u_{0,1}\|_{L^{2}}\leq C(1+\nu t)^{-\frac{\sigma}{2s}}\left(\|\Lambda_{1}^{-\sigma}\partial_{2}u_{0,1}\|_{L^{2}}+\|\partial_{2}u_{0,1}\|_{L^{2}}\right)\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma}{2s}}.$ (4.25)
@@
Similarly, for $0<\theta<\min\{\frac{1}{9},\frac{1-2\sigma}{2\sigma}\}$,

### v2

@@ 2 Anisotropic Estimates and Tools
If $\|u_{0}\|_{H^{k}}\leq\varepsilon$ with $\varepsilon$ sufficiently small, then a standard bootstrap argument yields
$\|u(t)\|_{H^{k}}^{2}+\nu\int_{0}^{t}\|\Lambda_{1}^{s}u(\tau)\|_{H^{k}}^{2}\,d\tau\leq C\varepsilon^{2}.$
Proof of Lemma 4.1.
We again split the estimate into the cases $1\leq\beta\leq k-1$ and $\beta=k$, namely
$\mathcal{N}_{\rm low}$ $:=\sum_{\beta=1}^{k-1}\int_{\mathbb{R}^{2}}\partial_{2}^{\beta}u\,\partial_{1}\partial_{2}^{k-\beta}u\,\partial_{2}^{k}u\,dx,$
$\mathcal{N}_{\rm top}$ $:=\int_{\mathbb{R}^{2}}\partial_{2}^{k}u\,\partial_{1}u\,\partial_{2}^{k}u\,dx.$
For the lower-order part $1\leq\beta\leq k-1$, by Hölder’s inequality,
$|\mathcal{N}_{\rm low}|\leq C\sum_{\beta=1}^{k-1}\|\Lambda_{1}^{s}\partial_{2}^{k-\beta}u\|_{L^{2}}\|\Lambda_{1}^{1-s}\big((\partial_{2}^{\beta}u)(\partial_{2}^{k}u)\big)\|_{L^{2}}.$
For the first factor, interpolation in the $x_{2}$ direction gives
$\|\Lambda_{1}^{s}\partial_{2}^{k-\beta}u\|_{L^{2}}\leq\|\Lambda_{1}^{s}u\|_{L^{2}}^{\frac{\beta}{k}}\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}^{1-\frac{\beta}{k}}.$
Using interpolation in the $x_{1}$ direction,
$\|\Lambda_{1}^{s}u\|_{L^{2}}\leq\|u\|_{L^{2}}^{1-s}\|\partial_{1}u\|_{L^{2}}^{s},$
hence
$\|\Lambda_{1}^{s}\partial_{2}^{k-\beta}u\|_{L^{2}}\leq\Big(\|u\|_{L^{2}}^{1-s}\|\partial_{1}u\|_{L^{2}}^{s}\Big)^{\frac{\beta}{k}}\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}^{1-\frac{\beta}{k}}.$
For the second factor, by Lemma 2.3,
$\|\Lambda_{1}^{1-s}[(\partial_{2}^{\beta}u)(\partial_{2}^{k}u)]\|_{L^{2}}\leq\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{\beta}u\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{k}u\|_{L^{2}}.$
By Gagliardo–Nirenberg inequality,
$\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{\beta}u\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\leq\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{\beta}u\|_{L^{2}}^{\frac{1}{2}}\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{\beta+1}u\|_{L^{2}}^{\frac{1}{2}}.$
Interpolating the $\partial_{2}$ derivatives between order 0 and $k$, we obtain
$\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{\beta}u\|_{L^{2}}$ $\leq\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}u\|_{L^{2}}^{1-\frac{\beta}{k}}\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{k}u\|_{L^{2}}^{\frac{\beta}{k}},$
$\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{\beta+1}u\|_{L^{2}}$ $\leq\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}u\|_{L^{2}}^{1-\frac{\beta+1}{k}}\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{k}u\|_{L^{2}}^{\frac{\beta+1}{k}}.$
Therefore,
$\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{\beta}u\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\leq\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}u\|_{L^{2}}^{1-\frac{2\beta+1}{2k}}\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{k}u\|_{L^{2}}^{\frac{2\beta+1}{2k}}.$
Since $\frac{3}{4}<s<\frac{11}{12}$, we have $\frac{7}{24}<\frac{3}{4}-\frac{s}{2}<\frac{3}{8}<1$, and thus
$\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}u\|_{L^{2}}\leq\|u\|_{L^{2}}^{1-(\frac{3}{4}-\frac{s}{2})}\|\partial_{1}u\|_{L^{2}}^{\frac{3}{4}-\frac{s}{2}}.$
Also,
$\|\Lambda_{1}^{\frac{3}{4}-\frac{s}{2}}\partial_{2}^{k}u\|_{L^{2}}\leq\|\partial_{2}^{k}u\|_{L^{2}}^{1-\theta_{1}}\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}^{\theta_{1}},\qquad\theta_{1}=\frac{3-2s}{4s}.$
Collecting the above estimates, we obtain —N_low—≤C ∥ u ∥_L^2^γ_1 ∥ ∂_1 u ∥_L^2^γ_2 ∥ ∂_2^k u ∥_L^2^γ_3 ∥ Λ_1^s ∂_2^k u ∥_L^2^γ_4, where γ_1 = β(1-s)k + ( 14 + s2 ) ( 1 - 2β+12k ), γ_2 = βsk + ( 34 - s2 ) ( 1 - 2β+12k ), γ_3 = 6s-34s ( 1 + 2β+12k ), γ_4 = 1 - βk + 3-2s4s ( 1 + 2β+12k ). By Young’s inequality, we further get
$|\mathcal{N}_{\rm low}|\leq{}$ $C\bigg[\Big(\|u\|_{L^{2}}^{\gamma_{1}}\|\partial_{1}u\|_{L^{2}}^{\gamma_{2}}\|\partial_{2}^{k}u\|_{L^{2}}^{\gamma_{3}-\gamma_{4}/2}\Big)^{\frac{2}{2-\gamma_{4}}}+C\|\partial_{2}^{k}u\|_{L^{2}}\,\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}^{2}\bigg].$
Combining the decay rates (4.1), (4.2) and (4.5), we obtain
$|\mathcal{N}_{\rm low}|\leq C(C_{0}\varepsilon)^{\frac{2(\gamma_{1}+\gamma_{2}+\gamma_{3}-\gamma_{4}/2)}{2-\gamma_{4}}}(1+t)^{-\Omega_{1}}+C\|\partial_{2}^{k}u\|_{L^{2}}\,\|\Lambda_{1}^{s}\partial_{2}^{k}u\|_{L^{2}}^{2},$
where Ω_1 = γ1σ+ γ2(σ+ 1)s(2-γ4). By direct computation, we get 1 - Ω_1 = 4k(4s - 3) - 4σ(2k - 1)12ks + 12βs - 6k - 6β+ 2s - 3. When $k>2$, the denominator is positive. Therefore, $1-\Omega_{1}<0$ iff the numerator is negative, namely 4k(4s - 3) - 4σ(2k - 1)¡0, which is equivalent to s¡34+σ2-σ4k. Since k¿σ1-2σ, and $s<\frac{1}{2}+\sigma$, we obtain s¡12+σ¡34+σ2-σ4k. Hence $\Omega_{1}>1$.
For the endpoint part $\beta=k$, denote it by $\mathcal{N}_{\rm top}$. By Hölder’s inequality and Sobolev embedding,
$|\mathcal{N}_{\rm top}|$ $=\left|\int_{\mathbb{R}^{2}}\big(\partial_{2}^{k}u\,\partial_{1}u\big)\,\partial_{2}^{k}u\,dx\right|$
$\leq C\|\partial_{1}u\|_{L^{2}_{x_{1}}L^{\infty}_{x_{2}}}\|\partial_{2}^{k}u\|_{L^{4}_{x_{1}}L^{2}_{x_{2}}}^{2}$
$\leq C\|\partial_{1}u\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}\partial_{2}u\|_{L^{2}}^{\frac{1}{2}}\|\Lambda_{1}^{\frac{1}{4}}\partial_{2}^{k}u\|_{L^{2}}^{2}.$
Using interpolation,
$\|\partial_{1}\partial_{2}u\|_{L^{2}}$ $\leq\|\partial_{1}u\|_{L^{2}}^{\frac{k-1}{k}}\|\partial_{1}\partial_{2}^{k}u\|_{L^{2}}^{\frac{1}{k}},$
$\|\Lambda_{1}^{\frac{1}{4}}\partial_{2}^{k}u\|_{L^{2}}^{2}$ $\leq C\|\partial_{2}^{k}u\|_{L^{2}}^{\frac{3}{2}}\|\partial_{1}\partial_{2}^{k}u\|_{L^{2}}^{\frac{1}{2}}.$
Hence,
$|\mathcal{N}_{\rm top}|\leq C\|\partial_{1}u\|_{L^{2}}^{\frac{1}{2}}\|\partial_{1}u\|_{L^{2}}^{\frac{k-1}{2k}}\|\partial_{1}\partial_{2}^{k}u\|_{L^{2}}^{\frac{1}{2k}}\|\partial_{2}^{k}u\|_{L^{2}}^{\frac{3}{2}}\|\partial_{1}\partial_{2}^{k}u\|_{L^{2}}^{\frac{1}{2}}.$
Applying Young’s inequality, we obtain
$|\mathcal{N}_{\rm top}|\leq{}$ $C\bigg[\left(\|\partial_{1}u\|_{L^{2}}^{\frac{2k-1}{2k}}\|\partial_{2}^{k}u\|_{L^{2}}^{\frac{5k-1}{4k}}\right)^{\frac{4k}{3k-1}}+C\|\partial_{2}^{k}u\|_{L^{2}}\|\partial_{1}\partial_{2}^{k}u\|_{L^{2}}^{2}\bigg].$
Combining the decay and boundedness estimates (4.1), (4.2), and (4.5), we get —N_top— ≤(C C_0ε)^9k-33k-1 (1+t)^-Ω_2 + C∥∂_2^k u∥_L^2 ∥∂_1 ∂_2^k u∥_L^2^2, where Ω_2 = σ+12s ×4k-23k-1. By $k>\frac{1}{1-2\sigma}$ and $s<\frac{1}{2}+\sigma$, we have $s<\frac{1}{2}+\sigma<\frac{\sigma+1}{2}\times\frac{4k-2}{3k-1}$ and thus $\Omega_{2}>1.$
@@ 4.2 The decay of $u_{1}$
$\|u_{1}(t)\|_{L^{2}}$ $\leq\|e^{\nu\Lambda_{1}^{2s}t}u_{1,0}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}(u\cdot\nabla u_{1})(\tau)\|_{L^{2}}\,d\tau$
@@
$\|e^{\nu\Lambda_{1}^{2s}t}u_{1,0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma}{2s}}\Big(\|\Lambda_{1}^{-\sigma}u_{1,0}\|_{L^{2}}+\|u_{0}\|_{L^{2}}\Big)$
@@ 4.3 The decay of $u_{2}$
$u_{2}(t)=e^{\nu\Lambda_{1}^{2s}t}u_{2,0}-\int_{0}^{t}e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)d\tau.$
@@
$\|u_{2}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}u_{2,0}\|_{L^{2}}+\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{1}\Delta^{-1}\nabla\times(u\cdot\nabla u)(\tau)\|_{L^{2}}\,d\tau.$
@@
$\|e^{\nu\Lambda_{1}^{2s}t}u_{2,0}\|_{L^{2}}$ $\leq C(1+\nu t)^{-\frac{\sigma+1}{2s}}\big(\|\Lambda_{1}^{-\sigma}\psi_{0}\|_{L^{2}}+\|u_{2,0}\|_{L^{2}}\big)$
@@
For $II_{1}$, by Lemma 2.7, we take out $\Lambda_{1}^{\sigma+\frac{2}{3}}$, and then use the fact that $\|\Lambda^{-\frac{1}{2}+\sigma}f\|_{L^{2}}\leq\|\Lambda_{1}^{-\frac{1}{2}+\sigma}f\|_{L^{2}}$ and applying Lemma 2.4, we have
@@
$\leq$ $\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}\Big[(\Lambda_{1}^{\frac{5}{6}}u_{1})u\Big]\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}\Big[u_{1}(\Lambda_{1}^{\frac{5}{6}}u)\Big]\|_{L^{2}}\Big)d\tau.$ (4.20)
@@
$\int_{0}^{t}(t-\tau)^{-\frac{\sigma+\frac{2}{3}}{2s}}\Big(\|\Lambda^{-\frac{1}{2}-\sigma}\Big[(\Lambda_{1}^{\frac{5}{6}}u_{1})u\Big]\|_{L^{2}}+\|\Lambda^{-\frac{1}{2}-\sigma}\Big[u_{1}(\Lambda_{1}^{\frac{5}{6}}u)\Big]\|_{L^{2}}\Big)d\tau$
@@ 4.4 The decay of $\partial_{2}u_{1}$
$\|\partial_{2}u_{1}(t)\|_{L^{2}}\leq\|e^{\nu\Lambda_{1}^{2s}t}\partial_{2}u_{1,0}\|_{L^{2}}+C\int_{0}^{t}\|e^{\nu\Lambda_{1}^{2s}(t-\tau)}\partial_{2}(u\cdot\nabla u)(\tau)\|_{L^{2}}d\tau.$
@@
$\|e^{\nu\Lambda_{1}^{2s}t}\partial_{2}u_{1,0}\|_{L^{2}}\leq C(1+\nu t)^{-\frac{\sigma}{2s}}\left(\|\Lambda_{1}^{-\sigma}\partial_{2}u_{1,0}\|_{L^{2}}+\|\partial_{2}u_{1,0}\|_{L^{2}}\right)\leq\frac{C_{0}}{4}\varepsilon(1+\nu t)^{-\frac{\sigma}{2s}}.$ (4.22)
@@
Similarly, we can also obtain a decay rate estimate of $\|\partial_{2}^{2}u_{1}(\tau)\|_{L^{2}}$ for $0<\theta<\min\{\frac{1}{9},\frac{1-2\sigma}{2\sigma}\}$ which will be used in the estimate for $III_{2}$,